OpenStudy (anonymous):

There are 1,660 megawatts of wind-generated electricity produced globally every year. This amount is equivalent to (1 point)1,660,000 watts. 1,660,000 kilowatts. 16,600,000 watts. 166,000 kilowatts.

OpenStudy (cj7529):

1660 Mega watts = 1660 x10^6 Watts = 1,660,000,000 =1,660,000Kilo Watts As, Kilo = 10^3 Mega = 10^6 Giga = 10^9 Tera = 10^12 .... cant remember any others but there ya go